Write a balanced equation for the double-replacement precipitation reaction described, using the smallest possible integer coefficients.

1. A precipitate forms when aqueous solutions of lead(II) nitrate and barium chloride are combined.

2. A precipitate forms when aqueous solutions of iron(III) sulfate and barium hydroxide are combined.

3. A reaction occurs when aqueous solutions of sodium hydroxide and hydrochloric acid are combined.

4. A reaction occurs when aqueous solutions of potassium hydroxide and hydrosulfuric acid are combined. Assume excess base.
Answer #1

Part a

when aqueous solutions of lead(II) nitrate Pb(NO3)2 and barium chloride(BaCl2) are combined and produces aqueous solution of barium nitrate Ba(NO3)2 and precipitate of lead chloride (PbCl2).

BaCl2(aq) + Pb(NO3)2(aq) = Ba(NO3)2(aq) + PbCl2(s)

Part b

when aqueous solutions of iron(III) sulfate Fe2(SO4)3

and barium hydroxide Ba(OH)2 are combined and produces aqueous solution of iron(III) hydroxide and precipitate of barium sulfate (BaSO4)

Fe2(SO4)3 (aq) + 3Ba(OH)2 (aq) = 3BaSO4(s) + 2Fe(OH)3(aq)

Part c

when aqueous solutions of sodium hydroxide (NaOH) and hydrochloric acid (HCl) are combined and produces aqueous solution of sodium chloride (NaCl) and water (H2O)

NaOH (aq) + HCl (aq) = NaCl (aq) + H2O (l)

Part d

when aqueous solutions of potassium hydroxide (KOH) and hydrosulfuric acid (H2SO4) are combined and produces potassium sulfate(K2SO4) and water

H2SO4(aq) + 2KOH(aq)→K2SO4(aq) + 2H2O(l ).